# Supercritical fluids as a distinct state of matter characterized by sub-short-range structural order

### v1

@@ Abstract
Figure 3: Transport coefficients. (a) Reduced viscosity $\eta^{*}=\eta\left(\sigma^{2}/\sqrt{\epsilon m}\right)$, where $\sigma$ and $\epsilon$ are length and energy parameters in the Lennard-Jones potential (see Methods), and $m$ is the atomic mass. The errors bars represent the standard error of the mean. (b) Reduced diffusion coefficient $D^{*}=D\sqrt{\left(\frac{m}{\epsilon\sigma^{2}}\right)}$. In the kinetic theory expression, $D_{0}=\frac{3}{8\rho\sigma^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$, we have used both the ideal gas form $\rho_{\rm ideal}(P,T)=k_{\rm B}P/T$ (ideal kinetic theory), and the exact $\rho(P,T)$ from simulation data (kinetic theory).
Radial distribution functions & physical length scales. We measure $g(r)$ using MD simulations at $T=195$ K, in the pressure range from 5 bar to 900 bar (see Fig. S1a). To provide a characterization of the SSR, we proceed with the simple method introduced above. Let us first focus on the second peak in $g(r)$ (see Fig. 2a). Based on the data of $g(r)$, a separation pressure $P_{2}\approx 55$ bar is determined: when $P>P_{2}$, the height of the second peak $h_{2}(P)$ increases with $P$; when $P<P_{2}$, the second peak cannot be identified within the given accuracy of the numerical data. The position of the second peak, $r_{2}\sim 1.95\sigma_{0}$, provides an estimate of the SSR length, i.e., $\xi(P=P_{2})=r_{2}$. Generalizing this analysis to the third and fourth peaks in $g(r)$, we obtain $\xi(P_{3})=r_{3}$ and $\xi(P_{4})=r_{4}$, where $P_{n}$ is the pressure below which the $n$-th peak in $g(r)$ cannot be identified, and $r_{n}$ is the position of the $n$-th peak (see Fig. S2). Unfortunately, the limitation of the current numerical precision does not allow us to perform such an analysis when $n>4$.
The data of $\xi(P)$, plotted in Fig. 2b, give a direct, though approximate, estimate of the SSR length scale in SCF. Below $P_{2}$, the $g(r)$ has only one peak, $g(r)\approx\exp[-V(r)/k_{\rm B}T]$, as given by the low-density theory Chandler (1987). In this case, $g(r)$ decays beyond the first peak without any oscillations – this is the definition of the gas state. On the other hand, $P_{2}$ should correspond to the gas-SCF boundary at $T=195$ K. Interestingly, $(P_{2}\approx 55\,\rm{bar},T=195\,\rm{K})$ coincides with the point $(P^{-}\approx 60\,\rm{bar},T=195\,\rm{K})$ on the $L^{-}$ line estimated independently in  Li and Jin (2024) using thermodynamics, confirming $L^{-}$ as the gas-SCF boundary. When $P>P_{2}$, a clear growth of the SSR length $\xi(P)$ is observed. If the numerical precision of $g(r)$ is sufficiently high, we expect that $\xi(P)$ would diverge around the liquid-SCF boundary. In sharp contrast, the length scale $r^{*}\equiv\rm{max}(r_{0},r_{1})$, obtained by exponential fitting of the large-$r$ $\tilde{h}(r)$ data (Fig. S3), is nearly a constant in the interested pressure range (Fig. 2b). The results show that $\xi$, which can be unambiguously distinguished from $r^{*}$, captures the onset of SSR in the SCF.
In the Supplementary Material, we verify the universality of these findings by considering a second isothermal cut crossing the $L^{\pm}$ lines (see Fig. S4).
@@
Transport across three states of matter. Next, we present clear signatures of the $L^{\pm}$ crossover lines in the transport coefficients, demonstrating how the emergence of SSR structural order influences their behavior. In the dilute limit, transport processes are well described by the well-known kinetic theory Chapman and Cowling (1990); Hansen and McDonald (2013). However, beyond the gas-SCF boundary, the SSR order provides non-negligible corrections to the transport coefficients - in particular, to the viscosity $\eta$, where the excess part $\eta_{\rm ex}$ that depends on the interaction potential and configurational structure becomes dominant compared to the kinetic part $\eta_{0}$ (see below).
@@
To analyze the simulation results and NIST data (see Fig. 3), we first examine the kinetic theory expressions valid in the dilute limit, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$ and $\eta_{0}=\frac{5}{16\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$ Chapman and Cowling (1990); Hansen and McDonald (2013). As shown in Fig. 3a, kinetic theory describes well the viscosity data below the $L^{-}$ line, where $\eta$ is a constant independent of pressure, as expected for gases. However, kinetic theory significantly fails above that line, as $\eta$ exhibits a rapid increase with pressure. The location where this happens is in very good agreement with the thermodynamic estimate of the $L^{-}$ crossover line Li and Jin (2024).
@@
To understand the origin of this deviation, one should note the approximations employed within such a formalism. Kinetic theory, as indicated by its name, only considers the kinetic effects, ignoring any explicit contributions from the interparticle potential. The theory also neglects the memory effects in particle motions by assuming that successive collisions are completely uncorrelated. Mathematically, this means that the memory function of the stress autocorrelation function $G(t)$ is given by $M_{G}(t)\sim\delta(t)$, where $\delta(t)$ is the delta function (see Supplementary Material). As we explicitly demonstrate, both approximations fail above the $L^{-}$ line. Moreover, it is important to emphasize that the microscopic mechanisms underlying fluid viscosity differ fundamentally between gases and liquids. In gases, viscosity arises from particle collisions, whereas in liquids it is governed by potential energy barriers, processes that kinetic theory does not account for. This supports our proposal that the $L^{-}$ line marks the crossover where potential effects become significant and the dynamics are no longer fully described by pair collisions, as in the gaseous regime.
@@
In general, the viscosity contains two terms, $\eta=\eta_{0}+\eta_{\rm ex}$, where $\eta_{0}$ is the kinetic contribution predicted by the kinetic theory, and $\eta_{\rm ex}$ comes purely from the interaction potential. Using this separation, the gas-SCF crossover line can be formally identified as the location at which $\eta_{\rm ex}\approx 0$. Lucks and Moser developed a formalism (hereafter referred to as LM theory) to compute $\eta_{\rm ex}$ based on a memory-function approach Lucas and Moser (1979a, b). In order to achieve a more quantitative description, we take the version of the LM theory that approximates the excess memory function by a simple exponential form $M_{G}(t)=ae^{-bt}$, with $a$ and $b$ explicitly related to $g(r)$ and $V(r)$ (see Supplementary Material). As $P$ increases above $L^{-}$, $\eta_{\rm ex}$ quickly exceeds $\eta_{0}$. Using $g(r)$ from simulations, the theoretical values of $\eta$ computed using the LM theory are very close to simulation and NIST data (see Fig. 3a). Surprisingly, LM theory works reasonably well for both $\eta$ and $D$ even above the $L^{+}$ crossover lines. It remains unclear whether this is merely a fortuitous coincidence. Finally, based on the data of $\eta$, we also define an effective length unit $\sigma_{\rm eff}$ which shows a qualitatively similar behavior of $\sigma_{0}/\sigma_{\rm eff}$ compared to $\xi/\sigma_{0}$ (see Fig. S5). This observation deserves further investigation.
@@
Indeed, $D_{0}$ computed using the full equation of state agrees well with simulation data even beyond the $L^{-}$ line, extending approximately up to the $L^{+}$ line. In contrast, if the ideal gas equation of state is used, i.e., $\rho_{\rm ideal}(P,T)=P/(k_{\rm B}T)$, the resulting “ideal kinetic theory” begins to fail around the $L^{-}$ line (see Fig. 3b). This further supports our interpretation of $L^{-}$ as the boundary where structural correlations beyond nearest neighbors become significant and the system departs from gas-like behavior. Moreover, this analysis highlights that the gas–SCF boundary can be identified through transport coefficients, whose behavior is influenced by the emergence of SSR structural order in the SCF phase.
@@
As already emphasized, above the liquid-SCF boundary (Frenkel-line), transport properties are dominated by a different microscopic mechanism. Meta-stable states in the potential landscape are formed, and the transitions between such states are realized by activation processes. In this regime, the Arrhenius law gives a relation between $D$ and the typical activation energy barrier $\Delta E$, $D\sim e^{-\Delta E/k_{\rm B}T}\sim e^{{\Delta S}/k_{\rm B}}$. The excess entropy $\Delta S$} can be approximately described by the two-particle entropy $S_{2}=-2\pi\rho k_{\rm B}\int_{0}^{\infty}{g(r)\ln[g(r)]-[g(r)-1]}r^{2}dr$ Dzugutov (1996), which can be directly computed from the $g(r)$ data. Leveraging on these arguments, we find that the scaling $D\sim e^{{S_{2}}/k_{\rm B}}$ deviates from the simulation data around $L^{+}$, as expected (see Fig. S6Demmel and Jakse (2025).
Figure 4: Velocity auto-correlation functions. (a) Fitting the tail of the VACF. Assuming a stretched exponential form, $Z(t)\sim e^{-(t/\tau)^{\beta}}$, the exponent $\beta$ is obtained by the slope of the linear fit in the log-log plot of $\ln[Z(t)]$ vs. $t$. (b) $\beta$ as a function of $P$. Error bars represent the standard deviation of the fitted values from ten independent fitting ranges.
@@
Velocity auto-correlation function (VACF). Besides the signatures of the SCF state in structure and transport, the gas-SCF boundary can be also identified by pure dynamic probes, namely the onset of non-trivial memory effects. We focus on the normalized VACF $Z(t)$ (Fig. S1b), which satisfies, $\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau$, where $M_{z}(t)$ is the memory function. Interestingly, the intermediate-time behavior of the VACF has been used to locate the liquid-SCF boundary (the Frenkel-2 line). In fact, at high pressures, the VACF displays a minimum after a fast decay; this minimum is gradually suppressed by decreasing pressure, which eventually disappears around the crossover pressure defined by the Frenkel-2 line Brazhkin et al. (2013).
@@
Here we propose that the large-time behavior of the VACF, but also the very short one incarnated in the so-called Einstein frequency, contain information that reflects the gas-SCF boundary. It can be shown that (see Supplementary Material), when the memory function is $M_{Z}\sim\delta(t)$ (memory-less) or exponential $M_{Z}(t)\sim e^{-t/\tau_{M}}$, the VACF decays exponentially at large times, $Z(t)\sim e^{-t/\tau}$. However, this simple exponential form is consistent with our simulation data only below the $L^{-}$ line. Above the $L^{-}$ line, we find that at late time (up to the precision of our numerical data) the VACF is well-approximated by a stretched exponential, $Z(t)=A\exp\left[\left(-\frac{t}{\tau}\right)^{\beta}\right]$, with $\beta<1$. In Fig. 4b we show the exponent $\beta$, obtained by fitting the data with the stretched exponential form (Fig. 4a), as a function of pressure. We find that the $L^{-}$ line coincides with the location at which $\beta$ approaches one. The theoretical origin of stretched exponential behavior of $Z(t)$ deserves to be investigated in the future.
@@
In addition, the signature of $L^{\pm}$ lines can also be found in the Einstein frequency $\Omega$ that appears in the very-short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$ Hansen and McDonald (2013) (see Fig. S7). These results imply that, beyond structural and transport properties, the dynamical behavior of the SCF state also clearly distinguishes it from both the gas and liquid phases, reinforcing its identification as an independent and distinct state of matter.
@@
Experimental signatures and proposals. Our findings provide a basis for experimental verification in future studies. Accurate measurements of $g(r)$ in the SCF state are becoming available using modern neutron scattering instruments Dove and Li (2022); Iwase et al. (2010) and the state-of-the-art data analysis tools Ma et al. (2020). This would allow one to directly extract the SSR length $\xi$ from the experimental data using the approach presented here. It is also possible to measure transport coefficients using neutron scattering techniques.
In fact, a recent experimental study has already reported accurate data of $D$ in supercritical methane along the $T=200$ K isotherm Ranieri et al. (2024). In Fig.5, we present the experimental data from Refs. Oosting and Trappeniers (1971); Ranieri et al. (2024), together with the $L^{\pm}$ lines determined based on the NIST EOS of methane. The experimental data are fully consistent with our theoretical expectation. The diffusion constant $D$ follows the ideal kinetic theory behavior $D\sim 1/P$ up to a pressure around the $L^{-}$ line as in Fig. 3b. Beyond the $L^{-}$ line a clear deviation from this scaling is observed. The SCF state between $L^{\pm}$ lines appears to coincide with a “transition regime” between the gaseous low-$P$ and the liquid high-$P$ behavior. In agreement with our simulation results, this observation provides a first experimental evidence of the two $L^{\pm}$ crossover lines and of the existence of the supercritical fluid state, clearly separated from the gas and liquid phases.
@@
Discussion. In this study, we demonstrate that the signatures of the boundary between gas and SCF states can be identified by various physical quantities, including structure descriptors such as the radial distribution function, transport coefficients, and the velocity auto-correlation function. There have been extensive studies on the other boundary, i.e., the liquid-SCF boundary. Our study completes the phase diagram above the critical point clarifying the existence of three independent states of matter characterized by distinct physical properties. (i) The liquid-SCF boundary is the limit beyond which the system does not behave like a typical liquid. Because the liquid state can be characterized by various features, ranging from thermodynamics to dynamics to structures, the deviation from the liquid behavior can be defined according to different criteria Fisher and Widom (1969); Brazhkin et al. (2012a, 2013, a). (ii) The gas-SCF boundary is the limit beyond which the system does not behave like a typical gas. This line is reflected in several physical properties and can be therefore defined using thermodynamics, structural order, transport coefficients and finally dynamical behavior. (iii) In-between is the SCF state that is neither liquid-like nor gas-like but rather a separate state of matter, whose properties cannot be easily rationalized as a simple linear superposition of gas and liquid features.
@@
It remains unclear whether our findings, and in particular the existence of a gas-SCF boundary, extend to systems with purely repulsive interactions, where no thermodynamic separation between liquid and gas phases, and no critical point, exist. So far, in these systems, only a boundary line between gas-like and liquid-like states has been discussed, e.g., Huang et al. (2023); Jiang et al. (2025).
@@
On the other hand, the proposed framework and analysis methods can be generalized to many other systems. In particular, it would be very interesting to study the nature of order in the quantum supercritical state Wang et al. (2024); Lv et al. (2024), in the context of quantum many-body correlations. Moreover, the dynamical signatures of the $L^{-}$ line can be explored by in the supercritical kinetics of black holes Zhao et al. (2025a); Li et al. (2025a, b) and holographic models as well Zhao et al. (2025b). Another future direction is to examine in more detail the analogy between the supercritical fluid state and the QCD phase diagram of nuclear matter. Interestingly, a three-state phase diagram presenting an intermediate partially deconfined phase, and naive similarities with our proposal depicted in Fig. 1(a), has been recently discussed Hanada and Watanabe (2023); Glozman et al. (2022); Fujimoto et al. (2025).
@@
In conclusion, returning to the fundamental question of “what separates a liquid from a gas” Brazhkin and Trachenko (2012), our revised answer points to the existence of a distinct supercritical fluid state, clearly distinguishable from both gas and liquid phases in terms of thermodynamic, structural, transport, and dynamical properties.

### v2

@@ Abstract
Figure 3: Transport coefficients. (a) Reduced viscosity $\eta^{*}=\eta\left(\sigma^{2}/\sqrt{\epsilon m}\right)$, where $\sigma$ and $\epsilon$ are length and energy parameters in the Lennard-Jones potential (see Methods), and $m$ is the atomic mass. The errors bars represent the standard error of the mean. (b) Reduced diffusion coefficient $D^{*}=D\sqrt{\left(\frac{m}{\epsilon\sigma^{2}}\right)}$. In the kinetic theory expression, $D_{0}=\frac{3}{8\rho\sigma^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$, we have used both the ideal gas form $\rho_{\rm ideal}(P,T)=k_{\rm B}P/T$ (ideal kinetic theory), and the exact $\rho(P,T)$ from simulation data (kinetic theory). (c) Reduced diffusion coefficient as a function of the two-particle structural entropy $S_{2}$. The blue dashed line represents an exponential fit of the data above the $L^{+}$ line, $D=1.18e^{S_{2}/k_{\rm B}}$.
More specifically, we measure $g(r)$ using MD simulations at $T=195$ K, in the pressure range from 5 bar to 900 bar (see Fig. S1a) and we focus on the second peak in $g(r)$ (see Fig. 2d). Based on the data of $g(r)$, a separation pressure $P_{2}\approx 55$ bar is determined: when $P>P_{2}$, the height of the second peak $h_{2}(P)$ increases with $P$; when $P<P_{2}$, the second peak cannot be identified within the given accuracy of the numerical data.
Below $P_{2}$, the $g(r)$ has only one peak, $g(r)\approx\exp[-V(r)/k_{\rm B}T]$, where $V(r)$ is the pair interparticle interaction, as given by the low-density theory Chandler (1987). In this case, $g(r)$ decays beyond the first peak without any oscillations – this is the definition of the gas state. On the other hand, $P_{2}$ should correspond to the gas-SCF boundary at $T=195$ K. Interestingly, $(P_{2}\approx 55\,{\rm bar}$, $T=195\,~\rm{K})$ coincides with the point $(P^{-}\approx 60\,{\rm bar}$, $T=195\,\rm{K})$ on the $L^{-}$ line estimated independently in  Li and Jin (2024) using thermodynamics, confirming $L^{-}$ as the gas-SCF boundary. In the Supplementary Material, we verify the universality of these findings by considering a second isothermal cut crossing the $L^{\pm}$ lines (see Fig. S2). This observation suggests that, from a structural perspective, the $L^{-}$ line corresponds to the disappearance of the second peak in $g(r)$.
Below, we reveal that the supercritical boundaries are also reflected in transport properties, specifically, the viscosity $\eta$ and the self-diffusion coefficient $D$, as well as in the dynamical behavior, such as the velocity autocorrelation function (VACF), $Z(t)$. These observations can also be derived from a liquid-state theory based on the memory-function approach and are supported by simulation results. The numerically estimated gas–SCF boundary based on these criteria aligns closely with the lower $L^{-}$ line independently identified in Ref. Li and Jin (2024) using a thermodynamic criterion, suggesting a clear physical significance for the latter.
@@
Kinetic features: transport coefficients. Next, we present clear signatures of the $L^{\pm}$ crossover lines in the transport coefficients. In the dilute limit, transport processes are well described by the well-known kinetic theory Chapman and Cowling (1990); Hansen and McDonald (2013). However, beyond the gas-SCF boundary, non-negligible corrections to the transport coefficients emerge - in particular, to the viscosity $\eta$, where the excess part $\eta_{\rm ex}$ that depends on the interaction potential and configurational structure becomes dominant compared to the kinetic part $\eta_{0}$.
Figure 4: Velocity auto-correlation functions. (a) Fitting the tail of the VACF. Assuming a stretched exponential form, $Z(t)\sim e^{-(t/\tau)^{\beta}}$, the exponent $\beta$ is obtained by the slope of the linear fit in the log-log plot of $\ln[Z(t)]$ vs. $t$. (b) $\beta$ as a function of $P$. Error bars represent the standard deviation of the fitted values from ten independent fitting ranges. (c) Einstein frequency $\Omega$ (Thz) as a function of pressure $P$. The power-law fits (dashed colored lines) highlight the two distinct behavior in the gas and liquid states.
@@
To analyze the simulation results and NIST data (see Fig. 3), we first examine the kinetic theory expressions, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$ and $\eta_{0}=\frac{5}{16\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$, which are valid in the dilute limit Chapman and Cowling (1990); Hansen and McDonald (2013). As shown in Fig. 3a, kinetic theory describes well the viscosity data below the $L^{-}$ line, where $\eta$ is a constant independent of pressure, as expected for gases. However, kinetic theory significantly fails above that line, as $\eta$ exhibits a rapid increase with pressure. The location where this happens is in very good agreement with the thermodynamic estimate of the $L^{-}$ crossover line Li and Jin (2024).
@@
To understand the origin of this deviation, one should note the approximations employed within such a formalism. Kinetic theory, as indicated by its name, only considers the kinetic effects, ignoring any explicit contributions from the interparticle potential. The theory also neglects the memory effects in particle motions by assuming that successive collisions are completely uncorrelated. Mathematically, this means that the memory function of the stress autocorrelation function $G(t)$ is given by $M_{G}(t)\sim\delta(t)$, where $\delta(t)$ is the delta function (see Supplementary Material Sec. S2). As we explicitly demonstrate, both approximations fail above the $L^{-}$ line. Moreover, it is important to emphasize that the microscopic mechanisms underlying fluid viscosity differ fundamentally between gases and liquids. In gases, viscosity arises from particle collisions, whereas in liquids it is governed by potential energy barriers, processes that kinetic theory does not account for. This supports our proposal that the $L^{-}$ line marks the crossover where potential effects become significant and the dynamics are no longer fully described by pair collisions, as in the gaseous regime.
@@
In general, the viscosity contains two terms, $\eta=\eta_{0}+\eta_{\rm ex}$, where $\eta_{0}$ is the kinetic contribution predicted by the kinetic theory, and $\eta_{\rm ex}$ comes purely from the interaction potential. Using this separation, the gas-SCF crossover line can be formally identified as the location at which $\eta_{\rm ex}\approx 0$ no longer holds. Lucks and Moser developed a formalism (hereafter referred to as LM theory) to compute $\eta_{\rm ex}$ based on a memory-function approach Lucas and Moser (1979a, b). In order to achieve a more quantitative description, we take the version of the LM theory that approximates the excess memory function by a simple exponential form $M_{G}(t)=ae^{-bt}$, with $a$ and $b$ explicitly related to $g(r)$ and $V(r)$ (see Supplementary Material Sec. S2). As $P$ increases above $L^{-}$, $\eta=\eta_{0}+\eta_{\rm ex}$ quickly exceeds $\eta_{0}$. Using $g(r)$ from simulations, the theoretical values of $\eta$ computed using the LM theory are very close to simulation and NIST data (see Fig. 3a). Surprisingly, LM theory works reasonably well for both $\eta$ and $D$ even above the $L^{+}$ crossover lines. It remains unclear whether this is merely a fortuitous coincidence.
@@
Indeed, $D_{0}$ computed using the full EOS agrees well with simulation data even beyond the $L^{-}$ line, extending approximately up to the $L^{+}$ line. In contrast, if the ideal gas equation of state is used, i.e., $\rho_{\rm ideal}(P,T)=P/(k_{\rm B}T)$, the resulting “ideal kinetic theory” begins to fail around the $L^{-}$ line (see Fig. 3b). This further supports our interpretation of $L^{-}$ as the boundary where structural correlations beyond nearest neighbors become significant and the system departs from gas-like behavior. Moreover, this analysis highlights that the gas–SCF boundary can be identified through transport coefficients, whose behavior is influenced by the emergence of SSR structural order in the SCF phase.
@@
As already emphasized, above the liquid-SCF boundary (Frenkel-line), transport properties are dominated by a different microscopic mechanism. Meta-stable states in the potential landscape are formed, and the transitions between such states are realized by activation processes. In this regime, the Arrhenius law gives a relation between $D$ and the typical activation energy barrier $\Delta E$, $D\sim e^{-\Delta E/k_{\rm B}T}\sim e^{{\Delta S}/k_{\rm B}}$. The excess entropy $\Delta S$} can be approximately described by the two-particle entropy $S_{2}=-2\pi\rho k_{\rm B}\int_{0}^{\infty}\{g(r)\ln[g(r)]-[g(r)-1]\}r^{2}dr$ Dzugutov (1996), which can be directly computed from the $g(r)$ data. Leveraging on these arguments, we find that the scaling $D\sim e^{{S_{2}}/k_{\rm B}}$ deviates from the simulation data around $L^{+}$, as expected (see Fig. 3c) Demmel and Jakse (2025).
@@
Dynamical fingerprints: velocity auto-correlation function (VACF). Besides the signatures of the SCF state in structural and transport properties, the supercritical boundaries can be also identified by pure dynamic probes. We focus on the normalized VACF $Z(t)$ (Fig. S1b), which satisfies, $\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau$, where $M_{z}(t)$ is the memory function. Interestingly, the intermediate-time behavior of the VACF has been used to locate the liquid-SCF boundary (the Frenkel-2 line). In fact, at high pressures, the VACF displays a minimum after a fast decay; this minimum is gradually suppressed by decreasing pressure, which eventually disappears around the crossover pressure defined by the Frenkel-2 line Brazhkin et al. (2013).
@@
Here we propose that the large-time behavior of the VACF, but also the very short one incarnated in the so-called Einstein frequency, contain information that reflects the gas-SCF boundary. It can be shown that (see Supplementary Material Sec. S2), when the memory function is $M_{Z}\sim\delta(t)$ (memory-less) or exponential $M_{Z}(t)\sim e^{-t/\tau_{M}}$, the VACF decays exponentially at large times, $Z(t)\sim e^{-t/\tau}$. However, this simple exponential form is consistent with our simulation data only below the $L^{-}$ line. Above the $L^{-}$ line, we find that at late time (up to the precision of our numerical data) the VACF is well-approximated by a stretched exponential, $Z(t)=A\exp\left[\left(-\frac{t}{\tau}\right)^{\beta}\right]$, with $\beta<1$. In Fig. 4b we show the exponent $\beta$, obtained by fitting the data with the stretched exponential form (Fig. 4a), as a function of pressure (see Supplementary Material Sec. S3 for additional details). We find that the $L^{-}$ line coincides with the location at which $\beta$ approaches one. The stretched exponential of VACF (with $\beta<1$) suggests the onset of non-trivial memory effects beyond $L^{-}$, but its theoretical derivation deserves to be investigated in the future.
@@
In addition, the signature of $L^{\pm}$ lines can also be found in the Einstein frequency $\Omega$ that appears in the very-short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$ Hansen and McDonald (2013) (see Fig. 4c). These results imply that, beyond structural and transport properties, the dynamical behavior of the SCF state also clearly distinguishes it from both the gas and liquid phases, reinforcing its identification as an intermediate state of matter.
@@
Experimental signatures and proposals. Our findings provide a basis for experimental verification in future studies. Accurate measurements of $g(r)$ in the SCF state are becoming available using modern neutron scattering instruments Dove and Li (2022); Iwase et al. (2010) and the state-of-the-art data analysis tools Ma et al. (2020).
It is also possible to measure transport coefficients using experimental techniques such as neutron scattering. In fact, a recent experimental study has already reported accurate data of $D$ in supercritical methane along the $T=200$ K isotherm Ranieri et al. (2024). In Fig. 5, we present the experimental data from Refs. Oosting and Trappeniers (1971); Ranieri et al. (2024), together with the $L^{\pm}$ lines determined based on the NIST EOS of methane. The experimental data are fully consistent with our theoretical expectation. The diffusion constant $D$ follows the ideal kinetic theory behavior $D\sim 1/P$ up to a pressure around the $L^{-}$ line as in Fig. 3b. Beyond the $L^{-}$ line a clear deviation from this scaling is observed. The SCF state between $L^{\pm}$ lines appears to coincide with a “transition regime” between the gaseous low-$P$ and the liquid high-$P$ behavior. In agreement with our simulation results, this observation provides a first experimental evidence of the two $L^{\pm}$ crossover lines and of the existence of the supercritical fluid state, clearly separated from the gas and liquid phases. During the finalization of our work, Ref. Khrapak et al. (2025) appeared, presenting a parallel theoretical analysis of the experimental data from Ref. Ranieri et al. (2024). At present, any connection between our results and the ideas put forward in Ref. Khrapak et al. (2025) remains unclear, but certainly warrants further investigation.
@@
Discussion. In this study, we demonstrate that the signatures of the supercritical fluid boundaries can be identified by various physical quantities, including structure descriptors such as the radial distribution function, transport coefficients, and the velocity auto-correlation function. Our study completes the phase diagram above the critical point clarifying the existence of three independent states of matter characterized by distinct physical properties. (i) The liquid-SCF boundary is the limit beyond which the system does not behave like a typical liquid. Because the liquid state can be characterized by various features, ranging from thermodynamics to dynamics to structures, the deviation from the liquid behavior can be defined according to different criteria Fisher and Widom (1969); Brazhkin et al. (2012a, 2013, a). (ii) The gas-SCF boundary is the limit beyond which the system does not behave like a typical gas. This line is reflected in several physical properties and can be therefore defined using thermodynamics, structural order, transport coefficients and finally dynamical behavior. (iii) In-between is the SCF state that is neither liquid-like nor gas-like but rather a separate state of matter, whose properties cannot be easily rationalized as a simple linear superposition of gas and liquid features. We find that, the Widom line defined by the maximum of the Ornstein-Zernike correlation length (Supplementary Material Sec. S4Xu et al. (2005), and the Widom delta identified according to the fractions of liquid-like and gas-like atoms (Supplementary Material Sec. S5Jun et al. (2018), both lie within the intermediate region bounded by $L^{\pm}$ lines.
@@
It remains unclear whether our findings, and in particular the existence of a gas-SCF boundary, extend to systems with purely repulsive interactions, where no thermodynamic separation between liquid and gas phases, and no critical point, exist. So far, in these systems, only a boundary line between gas-like and liquid-like states has been discussed, e.g., Huang et al. (2023); Jiang et al. (2025). Moreover, whether our results extend to more complex molecular fluids such as CO2 or H2O remains an open question and warrants further investigation.
@@
On the other hand, the proposed framework and analysis methods can be generalized to many other systems. In particular, it would be very interesting to study the nature of order in the quantum supercritical state Wang et al. (2024); Lv et al. (2024), in the context of quantum many-body correlations. Moreover, the dynamical signatures of the $L^{-}$ line can be explored in the supercritical kinetics of black holes Zhao et al. (2025a); Li et al. (2025a, b) and holographic models as well Zhao et al. (2025b). Another future direction is to examine in more detail the analogy between the supercritical fluid state and the QCD phase diagram of nuclear matter. Interestingly, a three-state phase diagram presenting an intermediate partially deconfined phase, and naive similarities with our proposal depicted in Fig. 1(a), has been recently discussed Hanada and Watanabe (2023); Glozman et al. (2022); Fujimoto et al. (2025).
@@
In conclusion, returning to the fundamental question of “what separates a liquid from a gas” Brazhkin and Trachenko (2012), our revised answer points to the existence of an intermediate supercritical fluid state, clearly distinguishable from both gas and liquid phases in terms of thermodynamic, structural, transport, and dynamical properties.